OpenStudy (anonymous):

Evaluate the expression (-3 -5)^2

OpenStudy (anonymous):

@lucaz

OpenStudy (lucaz):

it's (-3-5)^2? like (-3+(-5))^2?

OpenStudy (anonymous):

thas the answer too

OpenStudy (lucaz):

I mean what is inside the parantheses, just -3-5?

OpenStudy (anonymous):

ye its just that

OpenStudy (lucaz):

well, -3-5=-8 (-8)^2=(-8)*(-8)=64

OpenStudy (anonymous):

thank yo I ave one more if you can answer also

OpenStudy (lucaz):

what is it?

OpenStudy (anonymous):

Evaluate the expression, given that z = -5. -z^2 - z - 11

OpenStudy (lucaz):

substitute z by -5, what do you get?

OpenStudy (lucaz):

-(-5)^2-(-5)-11

OpenStudy (anonymous):

-25 - -5 -11

OpenStudy (lucaz):

yeah, add them

OpenStudy (anonymous):

is it -41

OpenStudy (lucaz):

-(-5) becomes +5

OpenStudy (lucaz):

it's -1*-5=5

OpenStudy (anonymous):

I mean -31

OpenStudy (lucaz):

right

OpenStudy (anonymous):

thank you dude